1. a supplier submits lots of fabric to a textile manufacturer. the customer wants to know if the lot average breaking strength exceeds 200 psi. if so, she wants to accept the lot. past experience indicates that a reasonable value for the variance of breaking strength is 100(psi)2. four specimens are randomly selected, and the average breaking strength observed is y=214 psi. is the customer accepting this lot at alpha=0.05?
1
Expert's answer
2021-06-16T09:32:48-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu=200"

"H_1:\mu>200"

This corresponds to a right-tailed test, for which a z-test for one mean, with known population standard deviation will be used.

Based on the information provided, the significance level is "\alpha=0.05," and the critical value for a right-tailed test is "z_c=1.6449."

The rejection region for this right-tailed test is "R=\{z: z>1.6449\}."

The z-statistic is computed as follows:


"z=\dfrac{\bar{x}-\mu}{\sigma/\sqrt{n}}"

"=\dfrac{214-200}{10/\sqrt{4}}=2.8"

Since it is observed that "z=2.8>1.6449=z_c," it is then concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu"  is greater than 200, at the "\alpha=0.05" significance level.


Using the P-value approach: The p-value is "p=P(z>2.8)=0.002555," and since "p=0.002555<0.05=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu" is greater than 200, at the "\alpha=0.05" significance level.


 Therefore, there is enough evidence to claim that the customer is accepting this lot at "\alpha=0.05."